Adult tickets for a play cost $14.00 and child tickets cost $3.00
If there were 26 people at a performance and the theater collected $199.00 from ticket sales, how many adults and how many children attended the play?

Answer by funmath(2933)

Let the number of adult tickets = x
The cost of adult tickets=14x
The rest of the 26 tickets are childrens, so let the amount of childrens tickets=26-x.
The cost of childrens tickets=3(26-x)
The problem you need to solve would be:
14x+3(26-x)=199
14x+78-3x=199
11x+78=199
11x+78-78=199-78
11x=121
11x/11=121/11
x=11
The number of adults is 11.
The number of children is 26-11=15.